# Importance-Weighted Non-IID Sampling for Flow Matching Models

### v1

@@ Abstract
Flow-matching models effectively represent complex distributions, yet estimating expectations of functions of their outputs remains challenging under limited sampling budgets. Independent sampling often yields high-variance estimates, especially when rare but with high-impact outcomes dominate the expectation. We propose an importance-weighted non-IID sampling framework that jointly draws multiple samples to cover diverse, salient regions of a flow’s distribution while maintaining unbiased estimation via estimated importance weights. To balance diversity and quality, we introduce a score-based regularization for the diversity mechanism, which uses the score function, i.e., the gradient of the log probability, to ensure samples are pushed apart within high-density regions of the data manifold, mitigating off-manifold drift. We further develop the first approach for importance weighting of non-IID flow samples by learning a residual velocity field that reproduces the marginal distribution of the non-IID samples. Empirically, our method produces diverse, high-quality samples and accurate estimates of both importance weights and expectations, advancing the reliable characterization of flow-matching model outputs. Our code will be publicly available on GitHub.
@@ 1 Introduction
Flow matching models [DBLP:conf/iclr/LipmanCBNL23] have become powerful tools for representing complex distributions [DBLP:conf/cvpr/GoPJKKK25, DBLP:conf/cvpr/LiKGLKKG25, DBLP:conf/cvpr/XingZHJHZLY25, DBLP:conf/cvpr/WangWZTLST25]. Beyond drawing individual samples, many applications require expectations of functions of model outputs [DBLP:journals/corr/abs-2502-03687]. Formally, for a fixed, pre-trained flow matching model with distribution $p(x)$ and a task-defined functional $f:\mathcal{X}\!\to\!\mathbb{R}^{d}$, we seek
@@
For example, in image inpainting, $X$ denotes a completed image, $f(X)$ may represent the probability of a target category produced by a classifier, and $\mu$ is the expected probability (under the same classifier) with respect to the generative model’s output distribution. A common practice is Monte Carlo with IID samples $X^{(1)},\dots,X^{(n)}\!\sim\!p$,
@@
However, sampling from modern flows is costly, so $n$ is small and $\hat{\mu}_{\text{IID}}$ often exhibits high variance. The issue is acute when $f(x)$ varies sharply across regions of the support (e.g., rare but with high-impact outcomes).
@@
To reduce variance under a fixed sampling budget, we draw the $n$ samples jointly, promoting coverage of distinct, salient regions of $p$. Intuitively, a diverse set can approximate $\mu$ more efficiently than independent draws. A two-outcome toy example illustrates the point: suppose $x\in\{a,b\}$ with probabilities $p_{a},p_{b}$ and $\mu=p_{a}f(a)+p_{b}f(b)$. With two IID samples, $\hat{\mu}_{\text{IID}}=\tfrac{1}{2}\big(f(X^{(1)})+f(X^{(2)})\big)$ and $\mathrm{Var}\big(\hat{\mu}_{\text{IID}}\big)=\tfrac{1}{2}\,p_{a}p_{b}\big(f(a)-f(b)\big)^{2}$. If we instead enforce diversity by drawing a pair that always contains both outcomes, the induced marginals are $p_{a}^{\prime}=p_{b}^{\prime}=\tfrac{1}{2}$. Using importance weights $w_{a}=p_{a}/p_{a}^{\prime}=2p_{a}$ and $w_{b}=2p_{b}$, the estimator $\hat{\mu}_{\text{N-IID}}=\tfrac{1}{2}\big(w_{a}f(a)+w_{b}f(b)\big)=\mu$ is exact and has zero variance. In realistic settings, rare events ($p_{b}\!\ll\!1$) with large |f(a)f(b)||f(a)-f(b)| can occur with risk-sensitive objectives (e.g. safety), so IID sampling can be unreliable even with many draws. This motivates joint, diversity-enhancing sampling paired with importance weighting.
@@
Let $X^{(1:n)}\!\sim\!p_{\text{joint}}$ denote $n$ samples drawn once from a joint distribution that encourages diversity. Denote by $p^{\prime}(x)$ the marginal density of any single element under $p_{\text{joint}}$. An unbiased estimator is then (see Appendix˜A for details)
@@
Recent work explores non-IID sampling to improve diversity. Particle Guidance [DBLP:conf/iclr/CorsoXBBJ24] couples multiple diffusion trajectories via a repulsive potential, while DiverseFlow [DBLP:conf/cvpr/MorshedB25] injects diversity into flow matching through a determinantal point process (DPP) objective [kulesza2012determinantal]. Both can be viewed as augmenting the sampling dynamics with a diversity velocity that pushes concurrent trajectories apart. Specifically, when jointly sampling $n$ trajectories $X^{(1:n)}_{t}$ from a flow-matching ODE, these methods can be formulated as solving $n$ coupled ODEs
@@
We propose Importance-Weighted Non-IID Sampling for Flow Matching Models, a framework that achieves both (G1) and (G2). An illustration appears in Figure˜1. For a pre-trained flow-matching model $v$ and $n$ coupled trajectories $\{X_{t}^{(i)}\}_{i=1}^{n}$, we constrain the direction of $g$ using the model’s score $\nabla_{x}\log p(x\,|\,t)$ so that diversity pushes samples within high-density regions (along the data manifold) rather than ejecting them off-manifold. Since $g$ will be normalized in practice, removing components that push samples off-manifold will amplify on-manifold components. This yields sets that are both diverse and high-quality, alleviating the core trade-off observed in prior work. To obtain unbiased estimates, we provide (to our knowledge) the first method to compute importance weights for non-IID, jointly sampled outputs from flow-matching models. Directly estimating the marginal $p^{\prime}(x)$ of the joint sampler is challenging because $X^{(1:n)}$ is drawn only once. We address this by learning a lightweight residual velocity $r_{\phi}(x,t)$ such that the perturbed flow
@@
matches the marginal distribution induced by the diversity-coupled sampler at t=1t\!=\!1. Since rectified-flow models [DBLP:conf/iclr/LiuG023] are widely used in recent flow-matching work [DBLP:conf/icml/EsserKBEMSLLSBP24, DBLP:journals/corr/abs-2506-15742], we focus on rectified flows here and include discussions for general flow-matching models in Sec.˜3.2. For rectified flows, learning such a residual velocity field is sufficient for importance-weight estimation, and a smaller network for $r_{\phi}$ than for $v$ can be used to reduce training and inference cost.
@@
We evaluate our method comprehensively. For accurate diagnosis, we first test on a Gaussian mixture model, where the true density (under IID sampling) is available in closed form. Our method improves sample diversity and quality and yields more accurate importance-weight estimates and semantic estimation. We then evaluate advanced conditional flow-matching models: Stable Diffusion 3.5 Medium [DBLP:conf/icml/EsserKBEMSLLSBP24] for text-to-image generation and FLUX.1-Fill-dev [DBLP:journals/corr/abs-2506-15742] for image inpainting. Although training residual velocity fields for these large models is infeasible due to computational cost, our method improves coverage of the output distribution as a plug-in. Across these tasks, our approach produces diverse, high-quality samples and accurate importance weight and expectation estimation, and demonstrates potential gains for large models.
Our contribution can be summarized as follows:
  • A non-IID sampling for flow matching. We propose a framework for joint, non-IID sampling in flow matching models, enabling efficient expectation estimation under fixed sampling budgets.
  • Score-based diversity regularization. We constrain diversity directions using model scores to maintain sample quality while enhancing coverage.
  • Importance weighting for joint samples. We introduce the first method to compute unbiased importance weights for non-IID flow samples.
2 Preliminaries and Related Work
@@ 2.1 Flow Matching Models
Flow matching models [DBLP:conf/iclr/LipmanCBNL23] produce samples by simulating an ordinary differential equation (ODE) from a simple base distribution to the target data distribution. Specifically, a sample $X_{1}$ is obtained by integrating
@@
where $n$ is the number of integration steps and $0\!=\!t_{0}\!<\!\cdots\!<\!t_{n}\!=\!1$. The evolution of the sample density $p_{t}(x)$ at a fixed $x$ is given by the continuity equation:
@@
One widely used variant of flow matching is the rectified flow model [DBLP:conf/iclr/LiuG023], which has been adopted by Stable Diffusion 3 [DBLP:conf/icml/EsserKBEMSLLSBP24] due to its effectiveness. In rectified flow, given a sample $X_{0}\sim p_{0}$ and a target sample $X_{1}$ (data), the intermediate state is defined by linear interpolation $X_{t}=(1-t)\,X_{0}+t\,X_{1}$. The velocity field is then trained by
@@
which leads to $v_{\theta}(x,t)\approx\mathbb{E}[\,X_{1}-X_{0}\mid X_{t}=x\,]$, meaning the learned velocity at any point $x$ approximates the expected remaining change needed to reach the target. An important consequence is that the model’s score function $s(x,t)$ can be computed in closed form from the velocity field and the current sample (see Appendix˜B for a derivation)
@@ 2.2 Enhancing Diversity of Flow Matching Models
When drawing multiple samples concurrently from a generative model, a key goal is to ensure they cover different high-probability regions of $p_{\theta}$—in other words, to encourage diversity. Recent methods for diffusion [Rombach_2022_CVPR] and flow [DBLP:conf/iclr/LipmanCBNL23] models have explored non-IID sampling strategies to increase diversity among $n$ joint samples [DBLP:conf/iclr/CorsoXBBJ24, DBLP:conf/cvpr/MorshedB25]. The core idea is to introduce a diversity objective $h(X^{(1:n)}_{t})$, where higher $h$ indicates a more diverse set.
@@
where $D_{ij}$ is the squared distance between samples $i$ and $j$ at time $t$, and $\text{med}(D)$ is the median of all $D_{ij}$ with $i\neq j$ (treated as a fixed constant to avoid affecting gradients). In practice, the diversity term can be computed using the predicted final samples $\hat{x}_{1}^{(i)}=x_{t}^{(i)}+v(x_{t}^{(i)},t)(1-t)$ [DBLP:conf/cvpr/MorshedB25]. We adopt this formulation in all experiments; we omit it from analytical discussion, as it does not affect the conclusions.
To increase $h$ during sampling dynamics, one can compute the gradient of $h$ with respect to each sample’s position as
@@
and use a scaled version as the diversity velocity $u$ (as in Equation˜4):
@@ 3 Methodology
Our approach addresses both (G1) diversity with quality and (G2) unbiasedness by modifying the sampling of a pre-trained flow. We generate a batch of $n$ samples jointly, $X^{(1:n)}=(X^{(1)},\dots,X^{(n)})$, from velocity $v(x,t)$ with base distribution $p_{0}(x)$. We encourage diversity while preserving on-manifold quality and assign an importance weight to each sample so that the expectation estimator remains unbiased (as in Equation˜3). We achieve this with two components: (1) score-based diversity velocity regularization that pushes samples apart primarily along high-density directions, and (2) a residual velocity for importance weighting added to $v$ to model each sample’s marginal distribution under the joint sampler.
@@ 3.1 Score-based Diversity Velocity Regularization
As in Sec.˜2.2, define a diversity objective $h\big(X^{(1:n)}_{t}\big)$ and compute
@@
A diversity-driven move $x_{t}\mapsto x_{t}+\delta x$ departs from the manifold if it reduces the log-density at time $t$, i.e., $\log p(x_{t}+\delta x,t)<\log p(x_{t},t)$. To avoid this, constrain the move to satisfy
@@
where $s(x,t):=\nabla_{x}\log p(x,t)$ is the score. Decompose
@@
For rectified flows, the score is computed directly from $v$ via Equation˜10. Since $v(x,t)$ is already computed at each step, applying score-based regularization incurs negligible additional overhead. After computing $g_{\mathrm{reg}}$ for all samples and jointly normalizing them, we obtain diversity velocities $u$ with amplified on-manifold components and reduced (or removed) off-manifold components.
@@
In Figure˜2, we show two joint samples $x_{t}^{1}$ and $x_{t}^{2}$. Since the scores satisfy $s_{t}^{1}\cdot g_{t}^{1}>0$ and $s_{t}^{2}\cdot g_{t}^{2}<0$, regularization is active only for the second sample. Thus, $(g_{t}^{1})_{\mathrm{reg}}=g_{t}^{1}$ and $(g_{t}^{2})_{\mathrm{reg}}=(g_{t}^{2})_{\perp}$ (the hard version). After normalization, the diversity velocities $u_{1},u_{2}$ push samples apart while avoiding departures from the manifold.
@@ 3.2 Importance Weight Estimation
matches the marginal distribution induced by the diversity-coupled sampler at $t=1$. Since rectified-flow models [DBLP:conf/iclr/LiuG023] are widely used [DBLP:conf/icml/EsserKBEMSLLSBP24, DBLP:journals/corr/abs-2506-15742], we focus on rectified flows, where learning such a residual velocity field is sufficient for importance-weight estimation, and a small network (compared to that for $v$) can be used to reduce training and inference cost.
Our method is also suitable for general flow matching models, but score functions then need to be learned, e.g. via score matching [DBLP:conf/iclr/0011SKKEP21]. We summarize the three ODEs in Definitions˜1, 2 and 3.
@@
$\mathrm{d}X_{t}=\big(v(X_{t},t)+r_{\phi}(X_{t},t)\big)\,\mathrm{d}t,\quad X_{0}\sim p_{0}.$ We denote its probability density as $p^{\prime\prime}_{\phi,t}(x)$. We train $r_{\phi}$ such that $p^{\prime\prime}_{\phi,1}(x)\approx p^{\prime}_{1}(x)$ at any $x$ generated by the joint ODEs.
@@
so that $w_{\phi,1}(x)\approx w(x)$ and $w_{\phi,0}(x)=1$. From Equation˜8, we derive the evolution of $\log w_{\phi,t}(x)$ at a fixed position $x$, as given in Theorem˜1. Since sampling proceeds along the coupled dynamics in Definition˜2,
@@
combining this with Theorem˜1 yields Theorem˜2, which characterizes how $w_{\phi,t}(X_{t}^{(i)})$ evolves along each trajectory. As discussed in Sec.˜3.3, estimating the density ratio along trajectories addresses key limitations of fixed-position estimation, and we therefore adopt this as our default method. For rectified flows, the computation simplifies further—requiring no additional score function estimation—as shown in Corollary˜1. All proofs are provided in Appendix˜C.
@@
Suppose $v$ and $v+r_{\phi}$ are rectified flows, then Theorem˜2 simplifies without learning additional score functions: $\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ $=\nabla_{x}\cdot r_{\phi}(X_{t}^{(i)},t)+\tfrac{tv(X^{(-i)}_{t},t)-X^{(-i)}_{t}}{1-t}\cdot r_{\phi}(X_{t}^{(i)},t)$ $+\tfrac{t}{1-t}\big(r_{\phi}(X_{t}^{(i)},t)-u\big(X^{(i)}_{t},X^{(-i)}_{t},t)\big)\cdot r_{\phi}(X_{t}^{(i)},t).$

### v2

# arXiv:2511.17812v2[cs.CV] 26 Feb 2026

@@ Abstract
Flow matching models effectively represent complex distributions, yet estimating expectations of functions of their outputs remains challenging under limited sampling budgets. Independent sampling often yields high-variance estimates, especially when rare but high-impact outcomes dominate the expectation. We propose a non-IID sampling framework that jointly draws multiple samples to cover diverse, salient regions of a flow matching model’s generative distribution. To balance diversity and quality, we introduce a score-based regularization for the diversity mechanism (SR), which uses the score function, i.e., the gradient of the log probability, to ensure samples are pushed apart within high-density regions of the data manifold, mitigating off-manifold drift. To enable unbiased estimation when desired, we further develop an approach for importance weighting of non-IID flow samples by learning a residual velocity field that reproduces the marginal distribution of the non-IID samples and by evolving importance weights along trajectories. Empirically, our method produces diverse, high-quality samples and accurate estimates of both importance weights and expectations, advancing the reliable characterization of flow matching model outputs. Our code will be publicly available on GitHub.
@@ 1 Introduction
Flow matching models [Lipman et al., 2023] are powerful tools for representing complex distributions. Beyond drawing individual samples, many applications require expectations of functions of model outputs [Favero et al., 2025]. Formally, for a fixed, pre-trained flow matching model with distribution $p(x)$ and a task-defined functional $f:\mathcal{X}\!\to\!\mathbb{R}^{d}$, we seek
@@
For example, in image inpainting, $X$ is a completed image, $f(X)$ could be the probability of a category produced by a classifier, and $\mu$ is its expectation under the generative distribution. A common practice is IID Monte Carlo with $X^{(1)},\dots,X^{(n)}\!\sim\!p$,
@@
However, sampling from modern flows is costly, so $n$ is small and $\hat{\mu}_{\text{IID}}$ can have high variance, especially for rare but high-impact outcomes.
@@
To reduce variance under a fixed sampling budget, we draw the $n$ samples jointly, promoting coverage of distinct, salient regions of $p$. Intuitively, a diverse set can approximate $\mu$ more efficiently than independent draws. A two-outcome toy example illustrates this: let $x\in\{a,b\}$ with probabilities $p_{a},p_{b}$ and $\mu=p_{a}f(a)+p_{b}f(b)$. With two IID samples, $\hat{\mu}_{\text{IID}}=\tfrac{1}{2}\big(f(X^{(1)})+f(X^{(2)})\big)$ and $\mathrm{Var}\big(\hat{\mu}_{\text{IID}}\big)=\tfrac{1}{2}\,p_{a}p_{b}\big(f(a)-f(b)\big)^{2}$. If we instead enforce diversity by drawing a pair that always contains both outcomes, the induced marginals are $p_{a}^{\prime}=p_{b}^{\prime}=\tfrac{1}{2}$. Using importance weights $w_{a}=p_{a}/p_{a}^{\prime}=2p_{a}$ and $w_{b}=2p_{b}$, the estimator $\hat{\mu}_{\text{N-IID}}=\tfrac{1}{2}\big(w_{a}f(a)+w_{b}f(b)\big)=\mu$ is exact and has zero variance. In realistic settings, some rare but non-negligible events (small $p_{b}$) can have high impact (large |f(a)f(b)||f(a)-f(b)|), so IID sampling can be unreliable even with many draws. This motivates joint, diversity-enhancing sampling paired with importance weighting.
@@
Let $X^{(1:n)}\!\sim\!p_{\text{joint}}$ denote $n$ samples drawn once from a joint distribution that encourages diversity. Denote by $p^{\prime}(x)$ the marginal density of any single element under $p_{\text{joint}}$. An unbiased estimator is then (see Appendix A for details)
@@
Recent work explores non-IID sampling to improve diversity. Particle Guidance [Corso et al., 2024] couples diffusion trajectories via a repulsive potential, while DiverseFlow [Morshed and Boddeti, 2025] injects diversity into flow matching via a determinantal point process objective [Kulesza et al., 2012]. Both can be viewed as augmenting the sampling dynamics with a diversity velocity that pushes concurrent trajectories apart. When jointly sampling $n$ trajectories $X^{(1:n)}_{t}$ from a flow matching model, these methods solve $n$ coupled ODEs
@@
We propose a score-regularized non-IID sampling framework with importance weights that achieves both (G1) and (G2) (Figure 1). For a pre-trained flow matching model $v$ and $n$ coupled trajectories $\{X_{t}^{(i)}\}_{i=1}^{n}$, we constrain the direction of $g$ using the model’s score $\nabla_{x}\log p(x\,|\,t)$ so diversity pushes samples within high-density regions (along the data manifold) rather than ejecting them off-manifold. Since $g$ will be normalized in practice, removing components that push samples off-manifold will amplify on-manifold components. This yields sets that are both diverse and high-quality, alleviating the core trade-off observed in prior work. To obtain unbiased estimates, we provide (to our knowledge) the first method to compute importance weights for non-IID, jointly sampled outputs from flow-matching models. Directly estimating the marginal $p^{\prime}(x)$ of the joint sampler is challenging because $X^{(1:n)}$ is drawn only once. We address this by learning a lightweight residual velocity $r_{\phi}(x,t)$ such that the perturbed flow
@@
matches the marginal distribution induced by the diversity-coupled sampler at t=1t\!=\!1. Since rectified-flow models [Liu et al., 2023] are widely used in recent flow matching work [Esser et al., 2024, Labs et al., 2025], we focus on rectified flows here and include discussions for general flow-matching models in Section 3.2. For rectified flows, learning such a residual velocity field is sufficient for importance-weight estimation, and a smaller network for $r_{\phi}$ than for $v$ can be used to reduce training and inference cost. Once trained for a given joint-sampling configuration, the residual can be reused across many test evaluations, so the training cost is amortized across the expectations and thus is negligible.
@@
We evaluate our method comprehensively. For accurate diagnosis, we first test on a Gaussian mixture model, where the true density (under IID sampling) is available in closed form. Our method improves sample diversity and quality and yields more accurate importance-weight estimates and expectation estimation. We then evaluate advanced conditional flow-matching models: Stable Diffusion 3.5 Medium [Esser et al., 2024] for text-to-image generation and FLUX.1-Fill-dev [Labs et al., 2025] for image inpainting, where our method improves coverage of the output distribution as a plug-in via score-based regularization. Across these tasks, our approach produces diverse, high-quality samples and accurate importance weight and expectation estimation, and demonstrates potential gains for large models.
Contribution and Significance. We proposed a non-IID sampling framework for flow matching. We constrain diversity directions using model scores to maintain sample quality while enhancing coverage. For unbiased expectation estimation, we further introduce a novel method to compute importance weights for non-IID flow samples. The correctness of the proposed method has been theoretically proven followed by comprehensive empirical validation of effectiveness. By open-sourcing our code, this work facilitates the management of the diversity–quality tradeoff of flow matching generative models and effective importance weight estimation for unbiased expectation estimation, all of which are fundamental and important in flow matching generative models.
2 Preliminaries & Related Work
@@ 2.1 Flow Matching Models
Flow matching models [Lipman et al., 2023] generate a sample $X_{1}$ by integrating an ODE from a base distribution to the target distribution
@@
where $0\!=\!t_{0}\!<\!\cdots\!<\!t_{N}\!=\!1$ and $N$ is the number of integration steps. The evolution of the sample density $p_{t}(x)$ at a fixed $x$ is given by the continuity equation:
@@
One widely used variant of flow matching is the rectified flow model [Liu et al., 2023], adopted by Stable Diffusion 3 [Esser et al., 2024] due to its effectiveness. In rectified flow, given a sample $X_{0}\sim p_{0}$ and a target sample $X_{1}$ (data), the intermediate state is defined by linear interpolation $X_{t}=(1-t)\,X_{0}+t\,X_{1}$. The velocity field is then trained by
@@
which leads to $v_{\theta}(x,t)\approx\mathbb{E}[\,X_{1}-X_{0}\mid X_{t}=x\,]$, meaning the learned velocity at any point $x$ approximates the expected remaining change needed to reach the target. An important consequence is that the model’s score function $s(x,t)$ can be computed in closed form from the velocity field and the current sample (Appendix B)
@@ 2.2 Enhancing Diversity of Flow Matching Models
When drawing multiple samples concurrently from a generative model, a key goal is to ensure they cover different high-probability regions of $p_{\theta}$—in other words, to encourage diversity. Recent methods for diffusion [Rombach et al., 2022] and flow [Lipman et al., 2023] models have explored non-IID sampling strategies to increase diversity among $n$ joint samples [Corso et al., 2024, Morshed and Boddeti, 2025]. The core idea is to introduce a diversity objective $h(X^{(1:n)}_{t})$, where higher $h$ indicates a more diverse set.
@@
where $D_{ij}$ is the squared distance between samples $i$ and $j$ at time $t$, and $\text{med}(D)$ is the median of all $D_{ij}$ with $i\neq j$ (treated as a fixed constant to avoid affecting gradients). In practice, the diversity term can be computed using the predicted final samples $\hat{x}_{1}^{(i)}=x_{t}^{(i)}+v(x_{t}^{(i)},t)(1-t)$ [Morshed and Boddeti, 2025]. We adopt this formulation in all experiments; we omit it from analytical discussion, as it does not affect the conclusions.
To increase $h$ during sampling dynamics, one can compute the gradient of $h$ with respect to each sample as
@@
and use a scaled version as the diversity velocity $u$ (as in Eq. (4)):
@@ 3 Methodology
Our approach addresses both (G1) diversity with quality and (G2) unbiasedness by modifying the sampling of a pre-trained flow. We generate a batch of $n$ samples jointly, $X^{(1:n)}=(X^{(1)},\dots,X^{(n)})$, from velocity $v(x,t)$ with base distribution $p_{0}(x)$. We encourage diversity while preserving on-manifold quality and assign an importance weight to each sample so that the expectation estimator remains unbiased (as in Eq. (3)). We achieve this with two components: (1) score-based diversity velocity regularization that pushes samples apart primarily along high-density directions, and (2) a residual velocity for importance weighting added to $v$ to model each sample’s marginal distribution under the joint sampler.
@@ 3.1 Score-based Diversity Velocity Regularization
As in Section 2.2, define a diversity objective $h\big(X^{(1:n)}_{t}\big)$ and compute
@@
A diversity-driven move $x_{t}\mapsto x_{t}+\delta x$ departs from the manifold if it reduces the log-density at time $t$, i.e., $\log p_{t}(x_{t}+\delta x)<\log p_{t}(x_{t})$. To avoid this, constrain the move to satisfy
@@
where $s(x,t):=\nabla_{x}\log p_{t}(x)$ is the score. Decompose
@@
For rectified flows, the score is computed directly from $v$ via Eq. (10). Since $v(x,t)$ is already computed at each step, applying score-based regularization incurs negligible additional overhead. After computing $g_{\mathrm{reg}}$ for all samples and jointly normalizing them, we obtain diversity velocities $u$ with amplified on-manifold components and reduced (or removed) off-manifold components.
@@
In Figure 2, we show two joint samples $x_{t}^{1}$ and $x_{t}^{2}$. Since the scores satisfy $s_{t}^{1}\cdot g_{t}^{1}>0$ and $s_{t}^{2}\cdot g_{t}^{2}<0$, regularization is active only for the second sample. Thus, $(g_{t}^{1})_{\mathrm{reg}}=g_{t}^{1}$ and $(g_{t}^{2})_{\mathrm{reg}}=(g_{t}^{2})_{\perp}$ (the hard version). After normalization, the diversity velocities $u_{1},u_{2}$ push samples apart while avoiding departures from the manifold.
@@ 3.2 Importance Weight Estimation
matches the marginal distribution induced by the diversity-coupled sampler at $t=1$. Since rectified-flow models [Liu et al., 2023] are widely used [Esser et al., 2024, Labs et al., 2025], we focus on rectified flows, where learning such a residual velocity field is sufficient for importance-weight estimation, and a small network (compared to that for $v$) can be used to reduce training and inference cost.
@@
$\dot{X}_{t}=v(X_{t},t)+r_{\phi}(X_{t},t),\quad X_{0}\sim p_{0}.$ We denote its probability density as $p^{\prime\prime}_{\phi,t}(x)$. We train $r_{\phi}$ such that $p^{\prime\prime}_{\phi,1}(x)\approx p^{\prime}_{1}(x)$ for $x$ generated by the joint ODEs.
Our method is also suitable for general flow matching models, but score functions then need to be learned, e.g., via score matching [Song et al., 2021]. We summarize the three ODEs in Definitions 1, 2, and 3.
@@
so that $w_{\phi,1}(x)\approx w(x)$ and $w_{\phi,0}(x)=1$. From Eq. (8), we derive the evolution of $\log w_{\phi,t}(x)$ at a fixed position $x$, as given in Theorem 1. Since sampling proceeds along the coupled dynamics in Definition 2,
@@
combining this with Theorem 1 yields Theorem 2, which characterizes how $w_{\phi,t}(X_{t}^{(i)})$ evolves along each trajectory. As discussed in Section 3.3, estimating the density ratio along trajectories addresses key limitations of fixed-position estimation, and we therefore adopt this as our default method. For rectified flows, the computation simplifies further—requiring no additional score function estimation—as shown in Corollary 1. All proofs are provided in Appendix C.
@@
Suppose $v$ and $v+r_{\phi}$ are rectified flows, then Theorem 2 simplifies without learning additional score functions: $\frac{\mathrm{d}}{\mathrm{dt}}\log w_{\phi,t}(X_{t}^{(i)})$ $=\nabla_{x}\cdot r_{\phi}(X_{t}^{(i)},t)+\tfrac{tv(X^{(i)}_{t},t)-X^{(i)}_{t}}{1-t}\cdot r_{\phi}(X_{t}^{(i)},t)$ $+\tfrac{t}{1-t}\big(r_{\phi}(X_{t}^{(i)},t)-u\big(X^{(i)}_{t},X^{(-i)}_{t},t)\big)\cdot r_{\phi}(X_{t}^{(i)},t).$